The following table shows some data for three zero-coupon bonds. The face value of each bond is $1,000. Yield to Maturity BonURGENT!!
Answer #1

Yield to maturity of zero coupon bond=(Face Value/Price)^(1/t)-1

1.
=(1000/220)^(1/20)-1
=7.865%

2.
=log(1000/220)/log(1+8%)
=19.67395075 years

3.
=1000/1.1^18
=179.8587899